OpenStudy (anonymous):

Find the derivative of f(x) = -12x2 + 9x at x = 6.

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@MrNood

OpenStudy (anonymous):

@ganeshie8

OpenStudy (mrnood):

\[12x ^{2}+9x\] Is that the correct formula?

OpenStudy (anonymous):

yes

OpenStudy (mrnood):

Sorry - I missed out the - sign at the beginning \[-12x ^{2}+9x\] Do you know how to get the derivative of \[y=-12x ^{2}\] and \[y=9x\]

OpenStudy (anonymous):

Well usually I would put it into the difference quotient formula, meaning i put (x+h) for every x in the original formula, then simplify, and then add the original formula to it , simplify that further and divide the entire thing by h.

OpenStudy (anonymous):

But when I originally did that, I got this (-33hx-33h-9x)/h which, I don't think is right

OpenStudy (anonymous):

So no I don't think I know how to do that. :C

OpenStudy (mrnood):

I am not familiar with the method you are describing do you know for instance that the derivitive of y=2x is 2 or that the derivative of \[y=x ^{2}\] is 2x

OpenStudy (anonymous):

No, my lesson didn't teach me that.. :(

ganeshie8 (ganeshie8):

whats the difference quotient formula you have been using ?

OpenStudy (anonymous):

These are my answer choices, -112.5 -135 -90 -108

OpenStudy (mrnood):

I'm sorry - the answer is not difficult - but I do not know the method that oyu are being taught. This a simple derivative and does not need any special formulae

OpenStudy (anonymous):

oh, the lesson only spoke about limits really

OpenStudy (anonymous):

could you teach me?

OpenStudy (anonymous):

And it is (limh->0) [f(x+h)-f(x)]/h

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

\[\large f'(x) = \lim \limits_{h\to 0}\dfrac{f(x+h) - f(x)}{h}\]

OpenStudy (anonymous):

yes but i am not famliiar with the ' symbol,

ganeshie8 (ganeshie8):

f(x) = -12x2 + 9x f(x+h) = ?

OpenStudy (anonymous):

-12(x+h)^2+9(x+h)

ganeshie8 (ganeshie8):

expand

OpenStudy (anonymous):

-12(x+h)(x+h)+9x+9h

OpenStudy (anonymous):

-12(x^2+2hx+2h)+9x-h

ganeshie8 (ganeshie8):

you should get : f(x+h) = -12(x^2+2hx+h^2)+9x+9h

OpenStudy (anonymous):

-24x^2-24hx-24h^2+9x+9h

ganeshie8 (ganeshie8):

f(x+h) = -12(x^2+2hx+h^2)+9x+9h = -12x^2 - 24hx -12h^2 + 9x + 9h

ganeshie8 (ganeshie8):

^^

ganeshie8 (ganeshie8):

next work the difference : f(x+h) - f(x)

ganeshie8 (ganeshie8):

f(x+h) - f(x) = -12x^2 - 24hx -12h^2 + 9x + 9h - (-12x^2 + 9x) = ?

OpenStudy (anonymous):

-24hx-12h^2+9h

ganeshie8 (ganeshie8):

excellent ! plug that value in your earlier difference quotient formula

OpenStudy (anonymous):

just over h or adding the original formula to it as well, because I thought we did that

OpenStudy (anonymous):

would it be, [-24hx-12h^2+9h]/h ?

ganeshie8 (ganeshie8):

thats the difference quotient, taking the limit gives u the derivative

ganeshie8 (ganeshie8):

\[\large f'(x) = \lim \limits_{h\to 0}\dfrac{-24hx -12h^2+9h}{h}\]

ganeshie8 (ganeshie8):

Notice that you can factor out `h` on numerator, and then cancel it with the denominator h

ganeshie8 (ganeshie8):

Do it.

OpenStudy (anonymous):

Oh sorry alright

OpenStudy (anonymous):

wait what hold on sorry

ganeshie8 (ganeshie8):

\[\large f'(x) = \lim \limits_{h\to 0}\dfrac{-24hx -12h^2+9h}{h}\] \[\large f'(x) = \lim \limits_{h\to 0}\dfrac{h(-24x -12h+9)}{h}\] \[\large f'(x) = \lim \limits_{h\to 0}-24x -12h+9\]

ganeshie8 (ganeshie8):

Now, you can take the limit because there is no h in the denominator

OpenStudy (anonymous):

OH

ganeshie8 (ganeshie8):

simply plugin h = 0 in the final expression

ganeshie8 (ganeshie8):

\[\large f'(x) = \lim \limits_{h\to 0}\dfrac{-24hx -12h^2+9h}{h}\] \[\large f'(x) = \lim \limits_{h\to 0}\dfrac{h(-24x -12h+9)}{h}\] \[\large f'(x) = \lim \limits_{h\to 0}-24x -12h+9\] \[\large f'(x) = -24x -12(0)+9\]

OpenStudy (anonymous):

WOW okay So -153?

ganeshie8 (ganeshie8):

\[\large f'(x) = -24x + 9\]

OpenStudy (anonymous):

I mean -135

ganeshie8 (ganeshie8):

Yep !

ganeshie8 (ganeshie8):

good job !!

OpenStudy (anonymous):

wow thank you so much I'm glad you helped me C:

ganeshie8 (ganeshie8):

np, you're welcome :)